# Improving the statistical efficiency of cross-conformal prediction

### v1

@@ Appendix E Additional experiments
The $\alpha$-level is set to $0.1$ and the conformal prediction methods are applied on $1000$ data points randomly sampled without replacement. The remaining part is used as a test set to compute the metrics. The procedure is repeated $20$ times to remove the randomness of the split and we report the averages over these $20$ trials. The methods used are cross-conformal prediction and its variants (with $K=10$), and split conformal prediction is added for comparison.
The results are reported in Figure 6, where it is possible to see that the smaller sets are obtained using the eu-mod-cross method. The modified variants using exchangeability and randomization exhibit higher variability in interval width, likely due to the use of randomization and the asymmetry of the combination rule. All proposed methods have an empirical coverage of at least $1-2\alpha$. We remark that cross-conformal prediction guarantees a coverage of at least $1-2\alpha$, but is usually conservative. The coverage of the new methods is closer to the target level $1-2\alpha$ and the new variants outperform cross-conformal prediction in terms of set size.
Figure 6: Results for the “Communities and Crime” dataset. Empirical size and empirical coverage of different conformal prediction algorithms are reported. The smaller sets are obtained using eu-mod-cross whose empirical coverage is around $0.85$. Cross-conformal prediction is conservative, but it tends to produce stable sets.
@@
We apply conformal prediction methods on a dataset of moderate dimensions, with $p=13$ and $n=506$. The aim is to predict the cost of a house in Boston given some information on the neighborhood. The algorithm used is standard linear regression. We apply conformal prediction methods using $200$ training points, the remaining part is used as test set. The number of different subsets for cross-conformal prediction is set to $K=5$ and the miscoverage rate is $\alpha=0.1$. The procedure is repeated $20$ times, and we report the averages over the $20$ replications.
From Table 3, we see that smaller sets are obtained on average using eu-mod-cross conformal prediction, while larger ones are produced by split conformal prediction, which exhibits high variability in set size. The methods e-mod-cross and u-mod-cross have an empirical coverage around $1-\alpha$, with an average size generally smaller than that obtained using cross-conformal prediction. Full conformal prediction exhibits low variability in terms of size, with the sets typically being smaller than those produced by split conformal prediction and cross conformal prediction. However, as already seen, these advantages are counterbalanced by a high computational cost.
@@
Mean 17.296 14.854 14.202 13.462 15.746 16.448 14.047
Sd 1.489 1.814 1.235 1.989 1.345 2.478 1.005
Median 17.086 14.351 14.168 12.968 15.704 16.725 14.213
Min 14.432 11.129 12.001 9.711 13.385 11.095 11.816
Max 20.800 18.559 16.894 17.590 19.101 21.562 15.408
Coverage 0.930 0.890 0.885 0.859 0.914 0.900 0.889
Table 3: Results for the “Boston Housing” dataset using OLS as regression algorithm. The results refer to set size, except for the last row, which refers to the marginal coverage. The smaller sets are obtained using eu-mod-cross conformal prediction. The variability is high when using split conformal prediction.
@@
We tested our methods on a dataset containing information on patients with early-stage Parkinson’s disease (Tsanas and Little,, 2009). The goal is to predict the total UPDRS (Unified Parkinson’s Disease Rating Scale) using a range of biomedical voice measurements. In particular, after some preprocessing operations, the data set includes $n=5875$ points and $p=13$ covariates. The two regression algorithms used are lasso regression (with penalty parameter equal to $0.01$) and random forest (with $25$ trees grown for each forest). The $\alpha$-level is set to $0.1$, the number of folds is $K=10$ and the conformal prediction methods are applied on $3000$ data points randomly sampled without replacement. The remaining part is used as a test set to compute the metrics. The procedure is repeated $20$ times to remove the randomness of the split. The results reported are the averages over these $20$ trials. We compare our proposals with cross-conformal prediction. In addition, split conformal prediction with miscoverage rate set to $\alpha$ and $2\alpha$ is added for comparison.
The results are reported in Table 4 and Table 5. In Table 4, we can see that for lasso regression the smaller sets are obtained using split conformal with miscoverage rate set to $2\alpha$. Overall, our approaches typically yield smaller sets compared to those obtained using standard cross-conformal prediction and split conformal prediction. However, we can observe a higher variability derived from the use of randomization (or sequential processing of the p-values). Interestingly, when the random forest is used as regression algorithm (Table 5), the smaller sets are obtained using the exchangeable and randomized cross-conformal prediction method. In particular, on average, the method also outperforms split conformal prediction trained at level $2\alpha$. In both cases, the marginal coverage of the proposed methods fluctuates between levels $1-2\alpha$ and $1-\alpha$.
@@
Mean 30.026 29.160 26.470 26.565 29.744 30.029 23.811
Sd 0.402 0.750 1.753 1.786 0.376 0.772 0.376
Median 30.056 29.175 26.312 26.533 29.725 30.333 23.801
Min 18.716 18.716 17.615 18.606 18.716 28.704 23.138
Max 32.808 30.826 30.936 30.716 32.698 31.355 24.883
Coverage 0.906 0.897 0.856 0.857 0.903 0.904 0.803
Table 4: Results for the UPDRS dataset using lasso as regression algorithm. The results refer to set size, except for the last row, which refers to the marginal coverage. On average the smaller sets are obtained using split conformal with miscoverage rate $2\alpha$. The marginal coverage of the proposed methods lies between $1-2\alpha$ and $1-\alpha$.
@@
Mean 17.236 15.368 14.806 13.940 17.073 19.429 14.623
Sd 0.915 1.101 1.482 1.408 0.905 0.716 0.751
Median 17.285 15.413 14.753 13.982 17.065 19.499 14.895
Min 8.808 7.376 7.156 5.725 8.697 18.035 13.246
Max 23.890 17.835 21.468 17.615 23.560 20.954 15.805
Coverage 0.931 0.884 0.887 0.847 0.929 0.901 0.806
Table 5: Results for the UPDRS dataset using random forest as regression algorithm. The results refer to set size, except for the last row, which refers to the marginal coverage. On average the smaller sets are obtained using the exchangeable and randomized version of cross-conformal. The marginal coverage of the proposed methods lies between $1-2\alpha$ and $1-\alpha$.
@@
The proposed methods are applied to the abalone dataset (Nash et al.,, 1994). The goal is to predict the age of abalones (the number of rings) using $p=8$ physical measurements. The dataset contains $n=4177$ observations where $4000$ observations are used as training points, while the remaining part is used as a test set. In this experiment, we directly modify the cross-conformal prediction as described in Section 4.4 (indeed, we remove the word mod from the labels in Table 6). The procedure is repeated $20$ times to remove the randomness of the split and the results reported are the average over the $20$ trials. The $\alpha$-level is set to $0.1$ and $K=10$. The regression algorithm used is a random forest with $25$ trees grown for each forest.
The results are reported in Table 6. The coverage level for the proposed method oscillates between levels $1-2\alpha$ and $1-\alpha$. The smaller sets are obtained on average by the split conformal prediction with a miscoverage rate equal to $2\alpha$. The suggested methods improve quite significantly the performance of cross-conformal prediction in terms of set size, although the variability is generally higher.
@@
Mean 6.858 6.342 5.671 5.512 6.848 4.806
Sd 0.041 0.251 0.059 0.280 0.217 0.096
Median 6.861 6.350 5.662 5.484 6.875 4.803
Min 6.750 5.704 5.586 4.900 6.351 4.595
Max 6.930 6.738 5.798 6.061 7.203 5.060
Coverage 0.901 0.877 0.855 0.836 0.895 0.795
Table 6: Results for the “Abalone dataset” using random forest as regression algorithm. The results refer to set size, except for the last row, which refers to the marginal coverage. The marginal coverage of the proposed methods lies between $1-2\alpha$ and $1-\alpha$.
@@ Appendix F Proofs of the results
Comparing Equation (6) with the set defined in Equation (8) we have that $\hat{C}^{\mathrm{mod\text{-}cross}}_{n,K,\alpha^{\prime}}(X_{n+1})$ coincide with $\hat{C}^{\mathrm{cross}}_{n,K,\alpha}(X_{n+1})$. The same result is obtained, for example, in Vovk et al., 2022a (, Chapter 4.4). The coverage statement and the properties regarding the size of the set are corollaries of Theorems 4.4, 4.6 and 4.7, applied with threshold $\alpha^{\prime}$. ∎

### v2

@@ Appendix E Additional experiments
The $\alpha$-level is set to $0.1$ and the conformal prediction methods are applied on $1000$ data points randomly sampled without replacement. The remaining part is used as a test set to compute the metrics. The procedure is repeated $100$ times to remove the randomness of the split and we report the averages over these $100$ trials. The methods used are cross-conformal prediction and its variants (with $K=10$), and split conformal prediction is added for comparison.
The results are reported in Figure 6, where it is possible to see that the smaller sets are obtained using the eu-mod-cross method. The modified variants using exchangeability and randomization exhibit higher variability in interval width, likely due to the use of randomization and the asymmetry of the combination rules. All proposed methods have an empirical coverage of at least $1-2\alpha$. We remark that cross-conformal prediction guarantees a coverage of at least $1-2\alpha$, but is usually conservative. The coverage of the new methods is closer to the level $1-2\alpha$ and the new variants outperform cross-conformal prediction in terms of set size.
Figure 6: Results for the “Communities and Crime” dataset. Empirical size and empirical coverage of different conformal prediction algorithms are reported. The $\alpha$-level is set to $0.1$. The smaller sets are obtained using eu-mod-cross whose empirical coverage is around $0.85$. Cross-conformal prediction is conservative, but it tends to produce stable sets.
@@
We apply conformal prediction methods on a dataset of moderate dimensions, with $p=13$ and $n=506$. The aim is to predict the cost of a house in Boston given some information on the neighborhood. The algorithm used is standard linear regression. We apply conformal prediction methods using $200$ training points, the remaining part is used as test set. The number of different subsets for cross-conformal prediction is set to $K=5$ and the miscoverage rate is $\alpha=0.1$. The procedure is repeated $100$ times, and we report the averages over the $100$ replications.
From Table 3, we see that smaller sets are obtained using eu-mod-cross conformal prediction, while larger ones are produced by split conformal prediction, which exhibits high variability in set size. The methods e-mod-cross and u-mod-cross have an empirical coverage slightly lower than $1-\alpha$, with an average size generally smaller than that obtained using cross-conformal prediction. Full conformal prediction exhibits low variability in terms of size, with the sets typically being smaller than those produced by split conformal prediction and cross-conformal prediction. However, as already seen, these advantages are counterbalanced by a high computational cost.
@@
Mean 17.303 14.920 14.143 13.370 15.753 16.357 14.516
Sd 1.440 1.998 1.152 1.899 1.307 2.751 1.244
Median 17.188 14.692 14.113 13.128 15.679 16.041 14.656
Min 13.883 10.068 11.551 8.822 12.766 11.730 11.998
Max 20.935 20.537 17.150 18.339 19.101 27.469 17.756
Coverage 0.927 0.888 0.878 0.855 0.908 0.897 0.888
Table 3: Results for the “Boston Housing” dataset using OLS as regression algorithm. The results refer to set size, except for the last row, which refers to the marginal coverage. The $\alpha$-level is set to $0.1$. The smaller sets are obtained using eu-mod-cross conformal prediction. The variability is especially high when using split conformal prediction.
@@
We tested our methods on a dataset containing information on patients with early-stage Parkinson’s disease (Tsanas and Little,, 2009). The goal is to predict the total UPDRS (Unified Parkinson’s Disease Rating Scale) using a range of biomedical voice measurements. In particular, after some preprocessing operations, the data set includes $n=5875$ points and $p=13$ covariates. The two regression algorithms used are lasso regression (with penalty parameter equal to $0.01$) and random forest (with $25$ trees grown for each forest). The $\alpha$-level is set to $0.1$, the number of folds is $K=10$ and the conformal prediction methods are applied on $3000$ data points randomly sampled without replacement. The remaining part is used as a test set to compute the metrics. The procedure is repeated $100$ times to remove the randomness of the split. The results reported are the averages over these $100$ trials. We compare our proposals with cross-conformal prediction. In addition, split conformal prediction with miscoverage rate set to $\alpha$ and $2\alpha$ is added for comparison.
The results are reported in Table 4 and Table 5. In Table 4, we can see that for lasso regression the smaller sets are obtained using split conformal with miscoverage rate set to $2\alpha$. Overall, our approaches typically yield smaller sets compared to those obtained using standard cross-conformal prediction and split conformal prediction. However, we can observe a higher variability derived from the use of randomization (or sequential processing of the p-values). Interestingly, when the random forest is used as a regression algorithm (Table 5), the smaller sets are obtained using the exchangeable and randomized cross-conformal prediction method. In particular, on average, the method also outperforms the split conformal prediction trained at level $2\alpha$. In both cases, the marginal coverage of the proposed methods fluctuates between levels $1-2\alpha$ and $1-\alpha$. On the other hand, from Table 5 it is possible to see that cross conformal is conservative with coverage higher than the nominal $1-\alpha$.
@@
Mean 29.965 28.956 26.427 26.252 29.686 29.901 23.773
Sd 0.404 1.019 1.744 1.942 0.383 0.878 0.402
Median 29.946 29.175 26.202 26.092 29.725 29.703 23.734
Min 11.450 10.019 8.477 9.248 11.340 28.539 22.913
Max 32.808 30.826 31.267 30.716 32.698 32.318 25.223
Coverage 0.904 0.892 0.854 0.850 0.901 0.901 0.800
Table 4: Results for the UPDRS dataset using lasso as regression algorithm. The results refer to set size, except for the last row, which refers to the marginal coverage. The $\alpha$-level is set to $0.1$. On average the smaller sets are obtained using split conformal with miscoverage rate $2\alpha$. The marginal coverage of the proposed methods lies between $1-2\alpha$ and $1-\alpha$.
@@
Mean 17.180 15.186 14.737 13.718 17.015 19.210 14.550
Sd 0.918 1.058 1.481 1.370 0.908 0.825 0.638
Median 17.175 15.303 14.643 13.762 16.954 19.232 14.541
Min 8.697 7.376 6.716 5.725 8.697 16.842 13.079
Max 23.890 17.945 22.129 17.725 23.560 20.954 16.551
Coverage 0.931 0.883 0.887 0.846 0.929 0.900 0.802
Table 5: Results for the UPDRS dataset using random forest as regression algorithm. The results refer to set size, except for the last row, which refers to the marginal coverage. The $\alpha$-level is set to $0.1$. On average the smaller sets are obtained using the exchangeable and randomized version of cross-conformal prediction. The marginal coverage of the proposed methods lies between $1-2\alpha$ and $1-\alpha$.
@@
The proposed methods are applied to the abalone dataset (Nash et al.,, 1994). The goal is to predict the age of abalones (the number of rings) using $p=8$ physical measurements. The dataset contains $n=4177$ observations where $4000$ observations are used as training points, while the remaining part is used as a test set. In this experiment, we directly modify the cross-conformal prediction as described in Section 4.4 (indeed, we remove the word mod from the labels in Tables 6, 7 and 8). The procedure is repeated $100$ times to remove the randomness of the split and the results reported are the average over the $100$ trials. The $\alpha$-level is set to $0.1$ and we use different number of folds, in particular, $K=\{5,10,20\}$. The regression algorithm used is a random forest with $25$ trees grown for each forest.
The results are reported in Tables 6, 7 and 8. The coverage level for the proposed method oscillates between levels $1-2\alpha$ and $1-\alpha$. The smaller sets are obtained on average by the split conformal prediction with a miscoverage rate equal to $2\alpha$. The suggested methods improve quite significantly the performance of cross-conformal prediction in terms of set size, although the variability is generally higher. There is a slight decrease in the size of sets, and a slight increase in variability, of the e and eu-mod-cross methods as $K$ increases. The results for the split conformal conformal prediction are essentially the same for all tables as a different number of folds is applicable just for cross conformal prediction and its variants.
@@
Mean 6.864 6.372 5.708 5.477 6.730 4.617
Sd 0.074 0.255 0.040 0.316 0.155 0.103
Median 6.858 6.332 5.713 5.390 6.703 4.583
Min 6.798 6.049 5.659 5.195 6.512 4.516
Max 6.986 6.762 5.755 6.015 6.912 4.777
Coverage 0.911 0.890 0.873 0.854 0.903 0.803
Table 6: Results for the “Abalone dataset” using random forest as regression algorithm. The results refer to set size, except for the last row, which refers to the marginal coverage. The number of folds is $K=5$ and the $\alpha$-level is set to $0.1$. The marginal coverage of the proposed methods lies between $1-2\alpha$ and $1-\alpha$.
cross e-cross u-cross eu-cross split split $(2\alpha)$
Mean 6.869 6.323 5.695 5.444 6.816 4.741
Sd 0.060 0.225 0.069 0.258 0.171 0.110
Median 6.865 6.360 5.692 5.450 6.815 4.750
Min 6.725 5.559 5.545 4.667 6.351 4.419
Max 7.030 6.800 5.878 6.061 7.235 5.060
Coverage 0.910 0.886 0.863 0.843 0.899 0.798
Table 7: Results for the “Abalone dataset” using random forest as regression algorithm. The results refer to set size, except for the last row, which refers to the marginal coverage. The number of folds is $K=10$ and the $\alpha$-level is set to $0.1$. The marginal coverage of the proposed methods lies between $1-2\alpha$ and $1-\alpha$.
cross e-cross u-cross eu-cross split split $(2\alpha)$
Mean 6.855 6.210 5.649 5.379 6.817 4.717
Sd 0.062 0.340 0.063 0.381 0.190 0.089
Median 6.850 6.283 5.642 5.403 6.837 4.728
Min 6.648 4.992 5.520 4.321 6.396 4.503
Max 7.032 6.760 5.787 6.315 7.358 4.971
Coverage 0.909 0.879 0.862 0.839 0.899 0.795
Table 8: Results for the “Abalone dataset” using random forest as regression algorithm. The results refer to set size, except for the last row, which refers to the marginal coverage. The number of folds is $K=20$ and the $\alpha$-level is set to $0.1$. The marginal coverage of the proposed methods lies between $1-2\alpha$ and $1-\alpha$.
Electricity consumption dataset.
We additionally analyze a real dataset on electricity consumption, where accurate uncertainty quantification is crucial as the supplier’s revenue depends on customer energy use. The dataset contains 35 4113541135\,41135 411 observations and $20$ covariates; 30 0003000030\,00030 000 observations are used for training, while the remaining ones are used for testing. The splitting is repeated $100$ times, and the algorithm used is a random forest (with ntree = 100). In this case, methods such as full conformal prediction or jackknife+ can be computationally expensive due to the relatively large number of observations, making sample-splitting-based methods preferable.
The results are presented in Table 9. The column “uneven split” represents split conformal prediction, where the training set comprises a $(K-1)/K$ fraction of the data points. This corresponds to split conformal prediction with the same fraction of training data points used by a single round of cross conformal prediction. The last two columns represent cases where the p-values are combined using the Bonferroni rule at level $\alpha$ (i.e., $K\min(\mathbf{p})$) or at level $2\alpha$ (i.e., $K/2\min(\mathbf{p})$) rather than the simple average. This implies that the corresponding prediction sets have coverage at least equal to $1-\alpha$ and $1-2\alpha$, respectively.
The smaller sets are obtained by the eu-mod-cross method, and in general there is an improvement in the set size if we use cross conformal prediction instead of split conformal prediction. Bonferroni method produces very large prediction sets and this is expected since the Bonferroni rule is not powerful when p-values (or sets) are highly dependent. Indeed, the Bonferroni correction is tightest when the p-values are nearly independent, while the conformal p-values across folds are highly dependent. On the other hand, the rule “twice the mean” used in cross conformal prediction is more powerful when p-values are dependent; see Section 6.1 Vovk and Wang, (2020) for a discussion. This aligns with Theorem 4 in Lei et al., (2018), which states that the use of multisplit conformal prediction in conjunction with the Bonferroni rule produces wide sets (specifically, under some assumptions, sets are wider than “single split” conformal prediction). The average set size of “uneven” split conformal prediction is smaller than that of split conformal prediction but slightly larger than that of cross conformal prediction. Moreover, its variability (Sd) is higher.
[…]
Mean 50.85 47.10 33.78 32.27 50.71 59.59 52.25 26.52 196.68 149.63
Sd 0.49 1.66 0.31 1.54 0.49 1.64 2.97 0.54 6.51 2.92
Median 50.83 47.38 33.80 32.47 50.69 59.64 52.12 26.47 197.46 149.71
Min 49.73 41.97 32.99 27.74 49.60 56.54 44.45 25.25 174.87 141.06
Max 52.38 49.67 34.41 36.20 52.25 64.95 59.84 27.96 210.94 157.85
Coverage 0.91 0.89 0.86 0.85 0.90 0.90 0.90 0.80 0.98 0.97
Table 9: Results for the “Electricity consumption dataset” using random forest as regression algorithm. The results refer to set size, except for the last row, which refers to the marginal coverage. The number of folds is $K=10$ and the $\alpha$-level is set to $0.1$. The marginal coverage of the proposed methods lies between $1-2\alpha$ and $1-\alpha$. The Bonferroni method gives large prediction sets.
@@ Appendix F Proofs of the results
Comparing Equation (6) with the set defined in Equation (8) we have that $\hat{C}^{\mathrm{mod\text{-}cross}}_{n,K,\alpha^{\prime}}(X_{n+1})$ coincides with $\hat{C}^{\mathrm{cross}}_{n,K,\alpha}(X_{n+1})$. The same result is obtained, for example, in Vovk et al., 2022a (, Chapter 4.4). The coverage statement and the properties regarding the size of the sets are corollaries of Theorems 4.4, 4.6 and 4.7, applied with threshold $\alpha^{\prime}$. ∎